# Price of Uncertainty for Consensus Games

### v1

@@ Abstract
Many game-theoretic models assume that players have access to accurate information, but uncertainty in observed data is frequently present in real-world settings. In this paper, we consider a model of uncertainty where adversarial perturbations of relative magnitude $1+\varepsilon$ are introduced to players’ observed costs. The effect of uncertainty on social cost is referred to as the price of uncertainty. We prove a tight bound on the price of uncertainty for consensus games of $\Theta(\varepsilon^{2}n^{2})$ for all $\varepsilon=\Omega\mathopen{}\left(n^{-1/4}\right)$. This improves a previous lower bound of $\Omega(\varepsilon^{3}n^{2})$ as well as a previous upper bound of $O(\varepsilon n^{2})$.
@@ 1 Introduction
If you are asked whether more of your friends like vanilla ice cream or chocolate ice cream, could you truly give the right answer with confidence? Even the best polling organizations sometimes fail to predict the outcomes of two-party majority elections. The truth is that uncertainty is everywhere and can impact our abilities to make rational decisions. However, most game-theoretic models assume that players have access to accurate information about interactions as well as their own costs. In the context of strategic interactions, it is interesting to see how uncertainty can affect agents’ behavior.
Balcan et al. (2009) consider how introducing small perturbations in observed costs in well-studied potential games can impact their behavior under improved-response and best-response dynamics. They coin the term price of uncertainty, and study it in a wide variety of potential games including set-covering games (Buchbinder et al., 2008), fair cost-sharing games (Anshelevich et al., 2008), job-scheduling games (Nisan et al., 2007), and consensus games. In some games, even extremely small perturbations can cause these dynamics to snowball out of control and move to states of much higher social cost, whereas in other games these dynamics are much more stable against large degrees of perturbation.
Consensus games are a special case of party affiliation games, which are in turn a class of potential games (Bhalgat et al., 2010). In a consensus game, each of $n$ players can choose between two parties: the white group and the red group. Players accrue a constant positive cost for interacting with those that chose the other side. Players are modeled as vertices of an undirected graph that can be colored either white or red, and a player’s cost is the number of neighbors with the other color, which we call bad neighbors. We will also call edges whose endpoints have different colors bad edges.
@@
Starting from a ground state with social cost at least 1, the price of uncertainty is the maximum multiplicative factor by which the social cost can increase via uncertain best-response dynamics. The term uncertain here refers to the fact that a player can miscount their cost, which recall is the number of bad neighbors, by a multiplicative factor of at most $1+\varepsilon$ for a small uncertainty factor $\varepsilon>0$ that can depend on $n$. Note that we can assume $\varepsilon=\Omega\left(\frac{1}{n}\right)$, otherwise a player cannot miscount its number of bad neighbors and hence there is no uncertainty. Balcan et al. (2011) prove an upper bound of $O(\varepsilon n^{2})$ on the price of uncertainty for consensus games that holds for all $\varepsilon=\Omega\left(\frac{1}{n}\right)$ as well as a lower bound of $\Omega(\varepsilon^{3}n^{2})$ that holds in the domain $\varepsilon=\Omega\left(n^{-1/3}\right)$.
@@
This paper obtains a lower bound for the price of uncertainty of consensus games of $\Omega\left(\varepsilon^{2}n^{2}\right)$ for all $\varepsilon=\tilde{\Omega}\left(n^{-1/2}\right)$. This is achieved using a construction that amplifies the snowball effect of uncertainty described by Balcan et al. (2011). Then, we will prove that the price of uncertainty for consensus games is also bounded by $O\left(\varepsilon^{2}n^{2}\right)$ for all $\varepsilon=\Omega\left(\varepsilon^{-1/4}\right)$ through the construction and analysis of a clever monovariant. These bounds coincide, so we conclude that the price of uncertainty for consensus games is $\Theta\left(\varepsilon^{2}n^{2}\right)$ for $\varepsilon=\Omega\left(n^{-1/4}\right)$, resolving an open problem proposed by Balcan et al. (2011).
@@ 2 Lower Bound
We first improve the lower bound for the price of uncertainty for consensus games. Furthermore, we do so in the expanded domain $\varepsilon=\tilde{\Omega}\left(n^{-1/2}\right)$. Here $\tilde{\Omega}$ hides factors that are logarithmic in $n$.
@@
The price of uncertainty for consensus games is $\Omega(\varepsilon^{2}n^{2})$ for $\varepsilon=\tilde{\Omega}\left(n^{-1/2}\right)$.
@@ 2.1 Proof of Improved Lower Bound
Define an uncertain best response as a best response that strictly increases the social cost as a result of uncertainty. By definition, an uncertain best response must involve a color switch. One general idea is to construct a graph as a sequence of levels where each level contains slightly more vertices than the previous level. Starting from the first level, there are enough bad edges for all vertices in the first level to play uncertain best responses, which increases the number of bad edges by a multiplicative factor of $1+\varepsilon$. The fact that the vertices in the first level switched color implies that the vertices in the second level now have enough bad neighbors to play an uncertain best response, and this effect snowballs across the levels.
@@ 3 Upper Bound
We first provide a simple proof, omitted in Balcan et al. (2011), that the price of uncertainty for consensus games is $O(\varepsilon n^{2})$.
Theorem 3 ((Balcan et al., 2011, Theorem 2)).
The price of uncertainty for consensus games is $O(\varepsilon n^{2})$.
Proof.
We prove that in order to increase the number of bad edges, there must have been $\Omega\left(\frac{1}{\varepsilon}\right)$ bad edges at the start. Consider the first step at which the number of bad edges increases. Note that in order for a vertex’s uncertain best response to increase the number of bad edges, the vertex must think it has more bad neighbors than it actually has, hence the number of bad edges must be $\Omega\left(\frac{1}{\varepsilon}\right)$. Since this is the first step at which the number of bad edges increases, the number of bad edges at the start must be $\Omega\left(\frac{1}{\varepsilon}\right)$. Since there can be at most $\binom{n}{2}=O(n^{2})$ bad edges at any step, we conclude that the price of uncertainty is $O(\varepsilon n^{2})$. ∎
3.1 Ideas for improving the upper bound
@@
The price of uncertainty for consensus games is $O(\varepsilon^{2}n^{2})$ for $\varepsilon=\Omega\left(n^{-1/4}\right)$.
Generally, the goal of an adversary that wants to maximally blow up the number of bad edges can be thought of as trying to maximize the number of additional bad edges that each new vertex produces in the uncertain best-response process. Consider a time in which there are $m^{2}$ bad edges. The most “compact” way to store these bad edges, that is, uses the least number of vertices, is a bipartite graph $K_{m,m}$. If each of the $m$ vertices on one side of the bipartite graph plays an uncertain best-response, we can blow up the number of bad edges by a multiplicative factor of $(1+\varepsilon)$. However, this requires $(1+\varepsilon)m$ additional new vertices in order to “absorb” the new bad edges produced by the uncertain best-response. Hence to increase the number of bad edges from $m^{2}$ to $(1+\varepsilon)m^{2}$ requires roughly $(1+\varepsilon)m$ new vertices.
For example, following this heuristic for constant $\varepsilon$, that is $\varepsilon$ that is not a function of $n$, the last stage will have a $K_{\Theta(\varepsilon n),\Theta(\varepsilon n)}$. Let us be optimistic in the sense that we assume that the original bad edges at each stage, which become good after vertices play an uncertain bad response, can be turned bad again. Assuming most edges can eventually be turned bad again as is the case in Section 2, the total number of bad edges at the end is
Since there must be $\Omega\left(\frac{1}{\varepsilon}\right)$ bad edges at the start as in Theorem 3, we estimate the upper bound on price of uncertainty as $O(\varepsilon^{2}n^{2})$.
Motivated by the intuition in Section 3.1, it appears natural to track the number of bad edges at every step and try to establish an upper bound on its growth rate as new vertices play uncertain bad responses. The hope is that we can prove an upper bound on the growth rate in line with the rate described in Section 3.1. Unfortunately, this approach fails, as illustrated by the following example.
Example 5.
Consider a white vertex $V$ connected to $m$ white vertices and $m+1$ red vertices for $m=\Omega\left(\frac{1}{\varepsilon}\right)$. The $m$ white vertices are all connected to another white vertex $W$ and the $m+1$ red vertices are all connected to another red vertex $R$. We claim that we can have $V$ play $2m+1$ best responses in order to roughly double the number of bad edges while only operating on these $O(m)$ vertices:
Note that Example 5 can be generalized to operate on a $K_{m,2m+1}$ where each vertex in the $m$-size part is white, $m$ of the vertices in the $(2m+1)$-size part is red, and $m+1$ of the vertices in the $(2m+1)$-size part is white to also roughly double the number of bad edges. The bad edge multiplicative growth rate of this process operating on $\Theta(m)$ vertices is 2, whereas in Section 3.1 it is $1+\varepsilon$. Hence using this analysis it is actually possible for there to be $\Omega(n^{2})$ bad edges at the last stage if operating on $m$ vertices can double the number of bad edges at each step. We thus cannot prove an upper bound of $O(\varepsilon n^{2})$ bad edges at the last stage using this strategy.
Instead, one clever idea is that we should focus not on the total number of bad edges at every step but rather consider the time step at which each vertex first plays an uncertain best response. This imposes a sequential order on the vertices from the first vertex to play an uncertain best response to the last. Recall that in order for a vertex with high degree to be able to play an uncertain best response, it must be the case that a large number of edges adjacent to this vertex must be bad. If we are to end up with $\Omega(\varepsilon n^{2})$ bad edges at the end, it must be the case that the later vertices in this sequence must have high degree on average. This requires a large number of existing bad edges to be generated from the vertices before the final vertices in the sequence, which in turn requires bad edges from the vertices before them, and so on.
@@ 3.2 Proof of Improved Upper Bound
We now give the formal proof of Theorem 4. It will be important to first establish a bound upon the starting number of bad edges.
Lemma 6.
In order to obtain a price of uncertainty on the order of $\Omega\left(\varepsilon^{2}n^{2}\right)$, the starting number of bad edges must be equal to $O\left(\varepsilon^{-2}\right)$.
@@
Let $G$ be the set of vertices of a graph upon which a consensus game is being played. Define $S_{0}$ as the set of vertices in $G$ that begin incident to at least one bad edge. Consider the set of vertices that begin without any bad edges and play at least one uncertain best response. They can be numbered in order of their first uncertain best response from first to last as $v_{1},v_{2},\dots,v_{m}$. Define a set $V$ as $\{v_{1},v_{2},\dots,v_{m}\}$. Finally, let $L$ be the set of vertices that begin without any bad edges and never play an uncertain best response. Notice that $S_{0}$, $V$, and $L$ are pairwise disjoint, and $S_{0}\cup V\cup L=G$.
For all integers $1\leq k\leq m$, let $S_{k}=S_{k-1}\cup\{v_{k}\}$.
@@
Let $k$ be an integer between 0 and m1m-1. Before $v_{k+1}$ plays its first uncertain best response, every bad edge is incident to at least one vertex in $S_{k}$. Furthermore, every bad edge is always incident to at least one vertex in $S_{m}$.
@@
Before any uncertain best responses are played, all bad edges are incident to vertices in $S_{0}$ by definition. Similarly, at this point, edges between two vertices in $G\setminus S_{0}$ must all be good.
Let $0\leq j\leq m-1$ be an integer. Suppose that Lemma 7 holds when $k=j$. Then, any edge between two vertices in $G\setminus S_{j}$ is good. By construction, the first vertex outside of $S_{j}$ to play an uncertain best response is $v_{j+1}$. When this happens, the only edges to be affected will be those incident to $v_{j+1}$. However, since $v_{j+1}\in S_{j+1}$, every bad edge must now be incident to at least one vertex in $S_{j+1}$. Finally, as long as no vertices outside of $S_{j+1}$ change colors, every bad edge will remain incident to at least one vertex in $S_{j}$. By induction, Lemma 7 must hold for $k$ between 0 and m1m-1.
Finally, by construction, all vertices in $G\setminus S_{m}$ never play an uncertain best response. This means that any edges incident to 2 of these vertices will remain good, so Lemma 7 holds for $k=m$ as well. ∎
Lemma 8.
If $\varepsilon=\omega\left(n^{-1/2}\right)$, it is optimal to have $m=\Theta(n)$.
Proof.
Clearly, $m=|V|\leq|G|=n$, so $m=O(n)$, so the upper bound holds. We will now attack the lower bound.
By construction, $G$ is comprised of the two disjoint sets $S_{m}$ and $L$. By Lemma 7, any vertices in $L$ that do not share an edge with a vertex in $S_{m}$ can never be incident to a bad edge. This means that we can delete these vertices along with all of their incident edges, which will always increase the final price of uncertainty. By construction, prior to when $v_{1}$ plays its first uncertain best response, any edges incident to 2 vertices in $G\setminus S_{0}$ must be good. Since $L=G\setminus S_{m}\subset G\setminus S_{0}$ and since vertices in $L$ never change their colors, any edges incident to two vertices in $L$ must always be good. By removing all such edges, we will always increase the price of uncertainty.
We are now left with a new set of vertices $L^{\prime}\subset L$ where every edge incident to a vertex in $L^{\prime}$ is also incident to a vertex in $S_{m}\setminus S_{0}=V$. For $1\leq k\leq m$, $v_{k}$ can only play its first uncertain best response if its number of bad edges is at least its number of good edges multiplied by $(1+\varepsilon)^{-1}$. By Lemma 7, the number of bad edges incident to $v_{k}$ is at most $|S_{k-1}|\leq|S_{m}|$. Right before $v_{k}$ plays its first uncertain best response, all edges incident to $v_{k}$ and a vertex in $L^{\prime}$ must be good, so
$|S_{m}|^{2}\cdot(1+\varepsilon)^{2}$ $\geq|S_{m}|\cdot|L^{\prime}|$
$|S_{m}|\cdot(1+\varepsilon)^{2}$ $\geq|L^{\prime}|$
$\Theta\left(|S_{m}|\right)$ $=\Theta\left(L^{\prime}\right).$
Thus, the number of vertices in the graph has the same asymptotic growth as that of $S_{m}$. Recall that $|S_{m}|=|S_{0}|+m$. By Lemma 6, we must have $O\left(\varepsilon^{-2}\right)$ edges that start as bad, and the number of vertices that begin incident to a bad edge is at most twice this quantity. So, $|S_{0}|=O\left(\varepsilon^{-2}\right)$ as well.

### v2

# The Price of Uncertainty for Social Consensus

@@ Abstract
How hard is it to achieve consensus in a social network under uncertainty? In this paper we model this problem as a social graph of agents where each vertex is initially colored red or blue. The goal of the agents is to achieve consensus, which is when the colors of all agents align. Agents attempt to do this locally through steps in which an agent changes their color to the color of the majority of their neighbors. In real life, agents may not know exactly how many of their neighbors are red or blue, which introduces uncertainty into this process. Modeling uncertainty as perturbations of relative magnitude $1+\varepsilon$ to these color neighbor counts, we show that even small values of $\varepsilon$ greatly hinder the ability to achieve consensus in a social network. We prove theoretically tight upper and lower bounds on the price of uncertainty, a metric defined by Balcan et al. (2009) to quantify the effect of uncertainty in network games.
@@ 1 Introduction
In the past decade the rapid expansion of social media platforms has resulted in the formation of large and highly connected social networks. The evolution of these networks has been extensively studied in order to better understand how opinions held by agents are affected by their interactions with other people. In a social setting, individuals generally tend to modify their beliefs to more closely align with others that they frequently interact with (Moussaïd et al., 2013; Garimella et al., 2018). This generally causes opinions held by agents in a network to generally converge over time as interacting agents attempt to reach a consensus. However, previous research has shown that additional factors, such as a sparse network structure or agents who refuse to change their opinion, can hinder consensus formation. (Acemoğlu et al., 2013; Hegselmann and Krause, 2002; Moussaïd et al., 2013). The vast majority of social network models in the existing literature assume that agents have accurate information about the opinions held by others. However, in large networks where vertices generally have many neighbors and where interactions between agents are often brief, agents are often unable to make accurate judgments about the opinions of other agents such as the number of their consenting or dissenting peers (Xiong and Liu, 2014).
Previous work modeling uncertainty in games has focused on the failure of agents to always pick the action with the optimal payoff. For example, in the quantal response model, agents sometimes deviate from optimal behavior with strategies providing higher payoffs being chosen more often than those providing lower payoffs (Penna, 2018; Blume, 1993; Alós-Ferrer and Netzer, 2010). Alternatively, agents may occasionally fail to participate in the game, which can sometimes increase the social cost (Meir et al., 2012; Li et al., 2017).
In this paper, we instead focus on the effect of uncertainty in agents’ observations rather than their actions. We model a social network as a graph with an agent at each vertex. Individuals have a choice between two colors, red or blue, and the goal of each agent is to be the same color as the majority of their neighbors. This is called a consensus game (Christodoulou et al., 2012; Fabrikant et al., 2004). We consider a model of uncertainty first proposed by Balcan et al. (2009) where agents experience small, adversarial perturbations of magnitude $1+\varepsilon$ to their observations of their neighbors. These perturbations capture the difficulty of ascertaining accurate information in large social networks, as the absolute margin of error increases proportionally to the number of neighbors an agent has. Alternatively, this model can also capture the effect of biased measurements, where agents may distort observations to agree with their beliefs (confirmation bias) or place undue emphasis on new information (novelty bias) (Ngampruetikorn and Stephens, 2016; Mao et al., 2024).
The goal of this paper is to study how uncertainty affects behavior in consensus games as agents make potentially suboptimal updates to their colors. Balcan et al. (2009) define a social cost function as the number of edges whose endpoints have different colors. They quantify the effect of uncertainty using a measure called the price of uncertainty (PoU) of consensus games, defined as the maximum multiplicative factor by which the social cost can increase from any starting state. Even extremely small perturbations can allow adversarial perturbations to snowball out of control and cause the game to move to states of much higher social cost (Balcan et al., 2011).
Our main results are tight asymptotic upper and lower bounds on the PoU for consensus games. Letting $n$ denote the number of agents participating in the consensus game, the previously best known asymptotic bounds due to Balcan et al. (2011) were a lower bound of $\Omega(n^{2}\varepsilon^{3})$ and an upper bound of $O(n^{2}\varepsilon)$. In this paper, we extend their work by proving that the PoU for consensus games is exactly $\Theta(n^{2}\varepsilon^{2})$. We believe that our result also holds irrespective of agent update ordering (i.e., the adversary can only control the magnitude of the perturbations), and this a potential future direction for research.
2 The Model
We consider a consensus game, where each of $n$ agents are located at distinct vertices of a simple undirected graph with $n$ vertices (Christodoulou et al., 2012). In this paper, we will use the words agent and vertex interchangeably. Each agent can choose a color from $\{0,1\}$. Let us refer to two agents as neighbors iff there exists an edge between the two agents. An agent’s cost is defined as the number of their neighbors that pick a different color, which we call bad neighbors. We will also call edges whose endpoints have different colors bad edges. Finally, for any state $S$ of the game, we define a social cost function $\operatorname{cost}(S)$ as the sum of the costs of every agent in the game.
@@ 2 The Model
Starting from a ground state $S_{0}$ with social cost at least 1, let $S_{t}$ be a state achieved after $t$ steps under uncertain best-response dynamics. Here, the term uncertain here refers to the fact that an agent might miscount their cost by a multiplicative factor of at most $1+\varepsilon$ for an uncertainty factor $\varepsilon>0$ that can depend on $n$. Define an uncertain best response as a best response that strictly increases the social cost as a result of uncertainty. By definition, an uncertain best response must involve a color switch. Note that we can assume $\varepsilon=\Omega\left(\frac{1}{n}\right)$, otherwise an agent cannot miscount its number of bad neighbors and hence there is no uncertainty. The price of uncertainty (PoU) of consensus games is defined as
$\operatorname{PoU}(\varepsilon,\text{consensus})=\max\left(\frac{\operatorname{cost}(S_{t})}{\operatorname{cost}(S_{0})}\right),$
where the maximum is taken over all initial states $S_{0}$ and all states $S_{t}$ reachable from $S_{0}$ in some number $t$ of steps (Balcan et al., 2009).
@@
This paper obtains a lower bound for the PoU of consensus games of $\Omega\left(\varepsilon^{2}n^{2}\right)$ for all $\varepsilon=\tilde{\Omega}(n^{-1/2})$. Here $\tilde{\Omega}$ hides factors that are logarithmic in $n$. This is achieved using an intricate construction that amplifies the “snowball” effect of uncertainty in Balcan et al. (2011). Next, we prove that the PoU for consensus games is also bounded from above by $O\left(\varepsilon^{2}n^{2}\right)$ for all $\varepsilon=\Omega(\varepsilon^{-1/4})$ by analyzing a clever monovariant. These bounds coincide, so we conclude that the asymptotically tight PoU for consensus games is $\Theta(\varepsilon^{2}n^{2})$ for $\varepsilon=\Omega(n^{-1/4})$, resolving an open problem of Balcan et al. (2011).
2.1 Main Technical Challenges
Lower Bound. In order to achieve a lower bound of $\Omega(\varepsilon^{3}n^{2})$, Balcan et al. (2011) utilizes an initializer gadget to set up initial colorings for an output component. This component consists of successive layers of vertices, where each layer has more vertices than the previous layer by a factor of $(1+\varepsilon)$, and every vertex is connected to all vertices on adjacent layers. This output component is responsible for generating $\Omega(\varepsilon n^{2})$ bad edges, and the initializer gadget requires $\Theta(1/\varepsilon^{2})$ starting bad edges, so this achieves a lower bound of $\Omega(\varepsilon^{3}n^{2})$. We have found that it is very difficult to increase the final number of bad edges produced by such an output for any construction because the graph immediately becomes too connected for enough agents to play uncertain best responses. Rather, we present an alternative construction for an initializer gadget in Section 3.1 that only requires $\Theta(1/\varepsilon)$ starting bad edges, which results in a construction for which the number of bad edges can increase by a factor of $O(\varepsilon^{2}n^{2})$.
Upper Bound. Generally, the goal of an adversary that wants to maximally blow up the number of bad edges can be thought of as trying to maximize the number of additional bad edges that each new vertex produces in the uncertain best-response process. Consider a time in which there are $m^{2}$ bad edges. The most “compact” way to store these bad edges, that is, uses the least number of vertices, is a bipartite graph $K_{m,m}$. If each of the $m$ vertices on one side of the bipartite graph plays an uncertain best-response, we can blow up the number of bad edges by a multiplicative factor of $(1+\varepsilon)$. However, this requires $(1+\varepsilon)m$ additional new vertices in order to “absorb” the new bad edges produced by the uncertain best-response. Hence to increase the number of bad edges from $m^{2}$ to $(1+\varepsilon)m^{2}$ requires roughly $(1+\varepsilon)m$ new vertices.
For example, following this heuristic for constant $\varepsilon$, that is $\varepsilon$ that is not a function of $n$, the last stage will have a $K_{\Theta(\varepsilon n),\Theta(\varepsilon n)}$. Let us be optimistic in the sense that we assume that the original bad edges at each stage, which become good after vertices play an uncertain bad response, can be turned bad again. Assuming most edges can eventually be turned bad again as is the case in our lower bound construction, the total number of bad edges at the end is
Since there must be $\Omega(1/\varepsilon)$ bad edges at the start, we estimate the upper bound on price of uncertainty as $O(\varepsilon^{2}n^{2})$.
It appears natural to track the number of bad edges at every step and try to establish an upper bound on its growth rate as new vertices play uncertain bad responses. The hope is that we can prove an upper bound on the growth rate in line with our estimate of $O(\varepsilon^{2}n^{2})$. Unfortunately, this approach fails, as illustrated by the following example.
Example 1.
Consider a white vertex $V$ connected to $m$ white vertices and $m+1$ red vertices for $m=\Omega(1/\varepsilon)$. The $m$ white vertices are all connected to another white vertex $W$ and the $m+1$ red vertices are all connected to another red vertex $R$. We claim that we can have $V$ play $2m+1$ best responses in order to roughly double the number of bad edges while only operating on these $O(m)$ vertices:
Note that Example 1 can be generalized to operate on a $K_{m,2m+1}$ where each vertex in the $m$-size part is white, $m$ of the vertices in the $(2m+1)$-size part is red, and $m+1$ of the vertices in the $(2m+1)$-size part is white to also roughly double the number of bad edges. The bad edge multiplicative growth rate of this process operating on $\Theta(m)$ vertices is 2, whereas in our initial estimate it was $1+\varepsilon$. Hence using this analysis it is actually possible for there to be $\Omega(n^{2})$ bad edges at the last stage if operating on $m$ vertices can double the number of bad edges at each step. We thus cannot prove an upper bound of $O(\varepsilon n^{2})$ bad edges at the last stage using this strategy.
Instead, one clever idea is that we should focus not on the total number of bad edges at every step but rather consider the step at which each vertex first plays an uncertain best response. This imposes a sequential order on the vertices from the first vertex to play an uncertain best response to the last. Recall that in order for a vertex with high degree to be able to play an uncertain best response, it must be the case that a large number of edges adjacent to this vertex must be bad. If we are to end up with $\Omega(\varepsilon n^{2})$ bad edges at the end, it must be the case that the later vertices in this sequence must have high degree on average. This requires a large number of existing bad edges to be generated from the vertices before the final vertices in the sequence, which in turn requires bad edges from the vertices before them, and so on.
@@ 3 Lower Bound
We first improve the lower bound for the PoU for consensus games. Furthermore, we do so in the expanded domain $\varepsilon=\tilde{\Omega}(n^{-1/2})$. Here $\tilde{\Omega}$ hides factors that are logarithmic in $n$.
@@
The PoU for consensus games is $\Omega(\varepsilon^{2}n^{2})$ for $\varepsilon=\tilde{\Omega}(n^{-1/2})$.
@@ 3.1 Proof of Improved Lower Bound
One general idea is to construct a graph as a sequence of levels where each level contains slightly more vertices than the previous level. Starting from the first level, there are enough bad edges for all vertices in the first level to play uncertain best responses, which increases the number of bad edges by a multiplicative factor of $1+\varepsilon$. The fact that the vertices in the first level switched color implies that the vertices in the second level now have enough bad neighbors to play an uncertain best response, and this effect snowballs across the levels.
@@ 4 Upper Bound
The PoU for consensus games is $O(\varepsilon^{2}n^{2})$ for $\varepsilon=\Omega(n^{-1/4})$.
In addition to focusing only on the steps at which each vertex plays its first uncertain best response, we will also carefully construct and analyze a monovariant that restricts the degrees of new vertices and allows us to bound the number of edges in the graph.
@@ 4.1 Proof of Improved Upper Bound
Let $G$ be a graph for a consensus game, and let $V$ be the set of vertices of $G$. It will be useful to first establish a bound upon the starting number of bad edges.
Proposition 5.
Suppose that it is possible to increase the number of bad edges in $G$ by a factor of $\Omega\left(\varepsilon^{2}n^{2}\right)$. Then, the starting number of bad edges must be $O\left(\varepsilon^{-2}\right)$.
@@
We begin by partitioning $V$ into three disjoint subsets:
  • Let $P$ be the set of vertices in $V$ that begin without any bad edges and play at least one best response.
  • Let $Q$ be the set of vertices in $V$ that begin without any bad edges and never play a best response.
  • Let $S_{0}$ be the set of all remaining vertices in $V$, i.e. those that begin incident to at least one bad edge.
Let us number the vertices in $P$ in order of their first best responses from first to last as $p_{1},p_{2},\dots,p_{m}$ for $m=|P|$. Then, for all integers $1\leq k\leq m$, define $S_{k}=S_{0}\cup\{p_{1},p_{2},\dots,p_{k}\}$. Now that we have set up our notation, we will prove some preliminary statements regarding the behavior of the consensus game.
Proposition 6.
Let $k$ be an integer such that $0\leq k\leq m$. Before $p_{k+1}$ plays its first best response, every bad edge is incident to at least one vertex in $S_{k}$. Furthermore, every bad edge is always incident to at least one vertex in $S_{m}$.
Proof.
Before any best responses are played, all bad edges have endpoints in $S_{0}$ by definition. Similarly, at this point, edges between two vertices in $V\setminus S_{0}$ must all be good.
Let $0\leq j\leq m-1$ be an integer. Suppose that Proposition 6 holds when $k=j$. Then, any edge between two vertices in $V\setminus S_{j}$ is good. By construction, the first vertex outside of $S_{j}$ to play a best response is $p_{j+1}$. When this happens, the only edges to be affected will be those incident to $p_{j+1}$. However, since $p_{j+1}\in S_{j+1}$, every bad edge must now be incident to at least one vertex in $S_{j+1}$. Finally, as long as no vertices outside of $S_{j+1}$ change colors, every bad edge will remain incident to at least one vertex in $S_{j}$. By strong induction, Proposition 6 must hold for all integers $k$ between 0 and m1m-1 inclusive.
Finally, by construction, all vertices in $V\setminus S_{m}=Q$ never play a best response. This means that any edges between 2 of these vertices will remain good, so Proposition 6 holds for $k=m$ as well. ∎
@@
Suppose that $\varepsilon=\omega(n^{-1/2})$. Then, the factor by which the number of bad edges in $G$ can be increased is maximized when $m=\Theta(n)$.
@@
Clearly, $m=|P|<|V|=n$, so $m=O(n)$ and the upper bound holds. We will now attack the lower bound.
By construction, $V$ consists of the two disjoint subsets $S_{m}$ and $Q$. By Proposition 6, any vertices in $Q$ that do not share an edge with a vertex in $S_{m}$ can never be incident to a bad edge. This means that we can delete these vertices along with all of their incident edges, which will never decrease the factor by which the number of bad edges in $G$ can be increased. By construction, prior to when $p_{1}$ plays its first best response, any edges between 2 vertices in $V\setminus S_{0}$ must be good. Since $Q=V\setminus S_{m}\subset V\setminus S_{0}$, and since vertices in $Q$ never change their colors, any edges between two vertices in $Q$ must always be good. By removing all such edges, we will never decrease the factor by which the number of bad edges in $G$ can be increased.
We are now left with a new set of vertices $Q^{\prime}\subset Q$ where every edge incident to a vertex in $Q^{\prime}$ is also incident to a vertex in $S_{m}\setminus S_{0}=P$. For integers $1\leq k\leq m$, $p_{k}$ can only play its first best response if the number of its bad neighbors is at least the number of its good neighbors multiplied by $(1+\varepsilon)^{-1}$. By Proposition 6, the number of bad edges incident to $p_{k}$ is at most $|S_{k-1}|<|S_{m}|$. Right before $p_{k}$ plays its first best response, all edges between $p_{k}$ and a vertex in $Q^{\prime}$ must be good, so
$|S_{m}|^{2}\cdot(1+\varepsilon)^{2}$ $\geq|S_{m}|\cdot|Q^{\prime}|$
$|S_{m}|\cdot(1+\varepsilon)^{2}$ $\geq|Q^{\prime}|$
$\Theta\left(|S_{m}|\right)$ $=\Theta\left(Q^{\prime}\right).$
Thus, the number of vertices in the graph has the same asymptotic growth as that of $S_{m}$. Recall that $|S_{m}|=|S_{0}|+m$. By Proposition 5, we must have $O\left(\varepsilon^{-2}\right)$ edges that start as bad, and the number of vertices that begin with a bad neighbor is at most twice this quantity. So, $|S_{0}|=O\left(\varepsilon^{-2}\right)$ as well.